OpenStudy (anonymous):

X= 8x+28 -------- 2(x+14) What of the following is a soultion for X? A.) x=7 B.)x=-2 C.)x=-14 D.)x=2

OpenStudy (anonymous):

yuki, u there?

OpenStudy (yuki):

yep

OpenStudy (anonymous):

can u help me on mine with the other parts?

OpenStudy (yuki):

katie, if you noticed that the right side can be simplified, the answer is easy

OpenStudy (anonymous):

??

OpenStudy (yuki):

azntiger: which problem?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

see them?

OpenStudy (anonymous):

wait can you help me first :)

OpenStudy (yuki):

of course, az, I will help you out in a couple of minutes

OpenStudy (yuki):

katie, which part are you having trouble ?

OpenStudy (anonymous):

no answer, none of the choices work

OpenStudy (anonymous):

okay thank

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

100%

OpenStudy (anonymous):

what answer would you chocie

OpenStudy (yuki):

katie, the answer is either -2 or 7 and let me explain

OpenStudy (anonymous):

yuki, can u help me after?

OpenStudy (yuki):

\[x = (8x+28)/2(x+14)\]

OpenStudy (yuki):

as: sure

OpenStudy (yuki):

so using cross multiplication... \[x*2(x+14) = 8x+28\]

OpenStudy (yuki):

once you simplify this you will get the form... \[x^2-5x-14 =0\]

OpenStudy (yuki):

do you think you can solve it from there katie?

OpenStudy (yuki):

let me know if you still need help

OpenStudy (yuki):

AZ, I'm answering your question

OpenStudy (radar):

When simplified, I come up with \[x^2+10x-14=0\]

OpenStudy (radar):

Providing answers of 1.25, and -11.25 Doesn't match any provided.